Four channels are multiplexed using TDM. If the input bit rate is 600 kbps and the multiplexer uses a time slot of 4 bits, what will be the frame duration and the bit duration in the output channel?
Answer #1

Each frame carries 4 bits from each channel;
the size of each frame, therefore, is 4 * 4 16 bits.
Because each channel is sending 600 kb/s and a frame carries 4 bits from each channel, the frame rate must be 600 kframes per second.
Frame Duration = 1/600,000 = 1.67µs
The bit rate is 600000 × 16, or 9600000 bps
Bit duration = 1/9600000 = 0.1 µs